How many pounds of cashews selling for $8.96 per pound must be mixed with 12 pounds of chocolates selling for $4.48 per pound to create a mixture that sells for $7.28 per pound?
Answer by mananth(16949)

How many pounds of cashews selling for $8.96 per pound must be mixed with 12 pounds of chocolates selling for $4.48 per pound to create a mixture that sells for $7.28 per pound?
cashews 8.96 / pound :: let it be x pounds
chocolates 4.48 /pound 12 pounds
mix 7.28 : x+12 pounds
...
8.96x+4.48*12=7.28*(x+12)
8.96x+53.76= 7.28x+87.36
8.96x-7.28x=87.36-53.76
1.68x=33.6
x = 20 pounds of cashews
...
m.ananth@hotmail.ca